OpenStudy (anonymous):

If A= (7.00 -9.00 ) units, B= (-9.00 + 4.00 ) units, and C= (28.0 + 20.0 ) units, determine a and b such that a A + b B +C = 0. a=? b=? (Large Letters are Vectors)

OpenStudy (anonymous):

Vectors have a distributive property. Do (7a, -9a) + (-9b, 4b) + (28, 20) = (0, 0) It's just algebra after that. In other words, 7a -9b + 28 = 0 -9a + 4b + 20 = 0 Solve for a and b and you're done! P.S. Strippers don't take physics.